OpenStudy (lordhelix8th):

Which circles have a circumference of 20 units? Choose all answers that are correct. A. a circle with a diameter of 40 units B. a circle with a radius of 10 units C. a circle with a diameter of 20 units D. a circle with a radius of 20 units

OpenStudy (lordhelix8th):

@iGreen

OpenStudy (igreen):

Circumference of ___ units?

OpenStudy (igreen):

Your question is incomplete.

OpenStudy (lordhelix8th):

sorry on top units=20

OpenStudy (igreen):

Okay. \(C = \pi d\) \(C = 2\pi r\)

OpenStudy (lordhelix8th):

so a and c

OpenStudy (igreen):

No..is that \(20 \pi\) units? In the question?

OpenStudy (lordhelix8th):

yes

OpenStudy (igreen):

A is incorrect..but C is correct.

OpenStudy (lordhelix8th):

b and c

OpenStudy (igreen):

\(C = \pi d\) \(C = \pi(20)\)

OpenStudy (igreen):

Correct